### v1

# Degree Matrix Comparison for Graph Alignment

Abstract
Graph alignment considers the optimal node correspondence across networks. To advance unsupervised graph alignment algorithms on plain graphs, we propose Degree Matrix Comparison (DMC). Through extensive experiments and mathematical motivations, we demonstrate the potential of this method. Remarkably, DMC achieves up to $99\%$ correct node alignment for $90\%$-overlap graphs and $100\%$ accuracy for isomorphic graphs. Additionally, we propose a reduced version of DMC (Greedy DMC) that provides a solution to the graph alignment problem with lower time complexity. DMC could significantly impact graph alignment, offering a reliable solution for the task.
1 Introduction
Graph alignment is a critical task with applications across various domains, like social networks, biology, and cybersecurity. It is also highly related to the subgraph isomorphism problem in mathematics and computer science. Hence, improving graph alignment methods holds significant value. Supervised or semi-supervised methods typically exploit node attributes, such as semantic features of users [1, 2], but these methods break down when users disguise their identities [3]. WAlign is a recent unsupervised method, but also harnesses the power of attributes [4]. We focus on the unsupervised graph alignment problem for unattributed plain graphs, leveraging geometric properties.
Our proposed method, Degree Matrix Comparison (DMC), is a method that fits the task. Unsupervised methods with a focus on graph geometry include: REGAL [5], FINAL [6], Klau [7], and IsoRank [8]. For a more comprehensive view of past unsupervised alignment methods, see [9]. DMC was constructed with two core ideas in mind: degree is the most accurate and direct description of local structure; moreover, in representing geometric structure, it makes sense to include both local and global information, like in [10, 11], especially in compact matrix form (adjacency matrices and Laplacian matrices).
We propose foregoing the need to know exactly which nodes are connected, like in an adjacency matrix, and record accurately the degree of a node and its neighbors’ degrees in a matrix. We show that DMC works well for aligning real world graphs with high clustering and link complexity.
2 Problem Set Up
To perform graph alignment, we generate a pair of graphs. There are two graph sampling methods. One of our methods utilizes random walk on networks to extract subgraphs [3, 12]. We take subgraph $G_{s}=(V_{s},E_{s})$, with $n=|V_{s}|$ distinct nodes, of a real world network $G_{r}=(V_{r},E_{r})$ through random walk. We pick random $v_{0}\in V_{r}$, and initiate continuous random walk on $G_{r}$ until we either get one connected component with $n$ distinct nodes, or exhaust a connected component in $G_{r}$ with insufficient nodes, whichever comes first. If process incomplete, we randomly pick a new node $v_{n}\in V_{r}$ and continue random walk until we obtain $V_{s}$ with $n$ distinct nodes. $E_{s}\subseteq E_{r}$ inherits edges between sampled nodes from $G_{r}$. To obtain $G_{1}=(V_{1},E_{1})\subseteq G_{s}$ and $G_{2}=(V_{2},E_{2})\subseteq G_{s}$, we perform random walk on $G_{s}$ to obtain a common region that gives percentage $p$ (with respect to $G_{s}$) of overlapping nodes between $G_{1}$ and $G_{2}$. Let $C\subseteq V_{s}$ denote the set of common nodes between $G_{1}$ and $G_{2}$, then $|C|=np$. We take $H_{1},H_{2}\subseteq(V_{s}\setminus C)$ such that $|H_{1}|=|H_{2}|$ and $V_{s}\setminus C=H_{1}\sqcup H_{2}$ (disjoint union). Then $V_{1}=C\sqcup H_{1}$, $V_{2}=C\sqcup H_{2}$, and $|V_{1}|=|V_{2}|$. $E_{1}\subseteq E_{s}$ is again the inherited set of edges, likewise for $E_{2}$.
In the other sampling method, $G_{1}=G_{s}=(V_{1},E_{1})$, and we delete edges of $G_{1}$ with probability $p_{d}$ to obtain another graph [5]. If $G_{2}=(V_{2},E_{2})$, then $V_{2}=V_{1}$, and $E_{2}\subseteq E_{1}$. If we have $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ for graph alignment, we try to find the optimal mapping $f:V_{1}\rightarrow V_{2}$ to maximize the number of $v_{1}\in C\subseteq V_{1}$ such that $v_{1}=v_{2}=f(v_{1})\in V_{2}$.
3 Proposed Method: Degree Matrix Comparison
Our proposed method requires creating two degree matrices and comparing them. We denote $\text{deg}(v)$ as the degree of a node $v$.
Definition 3.1 (Degree Matrices).
Consider a pair of graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$. By our problem set up, $N=|V_{1}|=|V_{2}|$. Let $V_{1}=\{v_{i}\}_{i=1}^{N}$ and $V_{2}=\{w_{j}\}_{j=1}^{N}$. Let $m_{1}=\text{max}(\{\text{deg}(v_{i})\}_{i=1}^{N})$ and $m_{2}=\text{max}(\{\text{deg}(w_{j})\}_{j=1}^{N})$, and $m=\text{max}(m_{1},m_{2})$. Then the degree matrix $M_{1}$ for $G_{1}$ is a matrix of dimension $(N,m)$, where each row left-aligns, for a unique $v_{1}\in V_{1}$, its neighbors’ degrees in ascending order, with zero-padding to the right. We define $M_{2}$ similarly.
The order of rows in $M_{1}$ and $M_{2}$ is arbitrary. After obtaining $M_{1}$ and $M_{2}$, we use the Munkres assignment algorithm to minimize cost between rows of the matrices, producing a map between rows of $M_{1}$ and $M_{2}$. We migrate this assignment directly to the assignment $f:V_{1}\rightarrow V_{2}$ between nodes. We call this entire process the Degree Matrix Comparison (DMC).
The complexity of DMC is at most $O(N^{3})$ for a pair of graphs that each has $N$ nodes [13, 14]. When $m<<N$, the complexity is $O(N^{2}\cdot m)<<O(N^{3})$. The complexity can be reduced through a combination of the auction algorithm and Munkres assignment algorithm, but accuracy is compromised [15]. The auction algorithm, with $O(N^{2})$ complexity and assignment threshold $\epsilon$, serves as a preparation step to reduce reliance on the Munkres assignment algorithm. We call this reduced version the Greedy DMC.
One theoretical motivation for the method is that degree matrices remain invariant in response to arbitrary node labeling. Moreover, a degree matrix ($N\times m$) encodes information more compactly than an adjacency matrix ($N\times N$). Additionally, combining local (degrees) and global (assignment) perspectives is important because the Friendship paradox shows how local observations can be completely shifted when we examine the same neighborhood globally [16, 17].
4 Experimental Results and Analysis
Here is a table that concludes the wide range of datasets we worked with.
COSNET [18] Flickr; Last.fm; Myspace
SNAP [19] Facebook; YouTube; CA, PA Roadmaps; PPI
NetworkRepository [20] All other networks.
Method Correct Percentage Approx. Complexity
DMC $96.09\%$ $\sim O(n^{2})-O(n^{3})$
Greedy DMC ($\epsilon=100$) $50.46\%$ $\sim O(n^{2})-O(n^{3})$
REGAL [5] $80\%<p<90\%$ $\sim O(n)-O(n^{2})$
FINAL [6] $35\%<p<45\%$ $\sim O(n^{2})$
Klau [7] $20\%<p<30\%$ $\sim O(n^{5})$
IsoRank [8] $5\%<p<15\%$ $\sim O(n^{4})$
Two-Step $4.42\%$ $\sim O(n^{2})-O(n^{3})$
Euc Dist. $0.15\%$ (near zero) $\sim O(n^{2})-O(n^{3})$
Table 1: Comparison with baselines on PPI network [4, 5, 6, 8]. Deletion probability $p_{d}=0.01$. “Two-Step” uses degrees for alignment, and refines it through average degrees of neighbors. “Euc Dist.” aligns two graphs by minimizing distances between embeddings. DMC has highest accuracy, but larger complexity compared to [5, 6].
(a) Social Networks
(b) Roadmaps
Figure 1: We use the random walk graph sampling method for these experiments.
Networks Mean Variance Total Nodes ($N$)
Flickr 52.36 5330.29 5000
Last.fm 26.03 2164.02 5000
Myspace 7.59 113.32 5000
Facebook 85.06 4040.27 700
YouTube 14.52 960.27 5000
CA Roadmap 2.55 0.96 1000
PA Roadmap 2.88 0.78 1000
Table 2: Means and variances of degree distributions of $G_{s}$. The best performing Facebook and YouTube have large variances in their degree distributions. Comparing Flickr, Last.fm, and Myspace, Flickr and Last.fm are much more heterogeneous. However, there is no strict correlation between variance and performance, as can be seen from California and Pennsylvania roadmaps.
Networks Percent Correct Density Avg. Clustering
bio-grid-fission-yeast .9932 .0123 .1874
bn-mouse-retina-1 .9900 .9983 1.5539
bio-ce-gn .9894 .0218 .1839
bn-fly-drosophila-medulla-1 .9378 .0211 3.8729
bio-grid-human .8440 .0014 .0800
bn-mouse-kasthuri .6156 .0032 0
bio-yeast .5822 .0018 .0708
Table 3: Comprehensive table of experiments on biological networks, ordered by DMC’s performance on $p=90\%$ overlap subgraphs. Density highlight threshold: .01, average clustering coefficient highlight threshold: .1. DMC works better on graphs with higher densities and clusterings.
When $\epsilon=0$, the Greedy DMC does not reduce to DMC, rather it is a different algorithm. This would be an algorithm that first finds exact matches of row vectors (same number of first neighbors and its neighbors’ degrees follow exact same distribution). The $\epsilon=0,10$ cases in Figure 2 (b) are worth exploring in the future due to their high performance scores. The methods still need to be studied since their exact time complexity can vary with respect to graphs with different structures. In certain cases, the algorithm can become very expensive. One could try to vary these two cases to strike a better balance between accuracy and time complexity.
(a) Performance Decay Plot
$\epsilon$ Correct Percentage
0 $96.58\%$
10 $96.32\%$
100 $50.46\%$
1000 $0.18\%$
(b) Greedy DMC Performance
Figure 2: Experiments on PPI, with random deletion sampling method.
We also applied DMC to graphs generated through the Erdos-Renyi (ER) and Barabasi-Albert (BA) models. We see that even for BA graphs, which are already closer to modeling real world networks, the DMC cannot achieve performance up to $90\%$. This relates to the active field of research on how to best depict real world networks mathematically.
(a) Discovery of negative linear relationship.
(b) BA model has much better performance than ER.
Figure 3: These results support previous analysis that DMC works well for heterogeneous graphs.
5 Conclusion and Future Work
We proposed the DMC and Greedy DMC, showed that they were theoretically motivated, and analyzed experimental results. DMC’s accuracy on classic PPI networks proves to be higher than carefully chosen baselines, and it works best for heterogeneous graphs. Future works could examine special cases of Greedy DMC, as pointed out before.
6 Acknowledgements
We thank Allison Gunby-Mann for introducing Ashley to the graph alignment problem and bringing the random walk sampling method to attention.