### v2

# Degree Matrix Comparison for Graph Alignment

Abstract
Graph alignment considers the optimal node correspondence across networks. To advance unsupervised graph alignment algorithms on plain graphs, we propose Degree Matrix Comparison (DMC). Through extensive experiments and mathematical motivations, we demonstrate the potential of this method. Remarkably, DMC achieves up to $99\%$ correct node alignment for $90\%$-overlap graphs and $100\%$ accuracy for isomorphic graphs. Additionally, we propose a reduced version of DMC (Greedy DMC) that provides a solution to the graph alignment problem with lower time complexity. DMC could significantly impact graph alignment, offering a reliable solution for the task.
1 Introduction
Graph alignment is a critical task with applications across various domains, like social networks, biology, and cybersecurity. It is also highly related to the subgraph isomorphism problem in mathematics and computer science. Hence, improving graph alignment methods holds significant value. Supervised or semi-supervised methods typically exploit node attributes, such as semantic features of users [1, 2], but these methods break down when users disguise their identities [3]. WAlign is a recent unsupervised method, but also harnesses the power of attributes [4]. We focus on the unsupervised graph alignment problem for unattributed plain graphs, leveraging geometric properties.
Our proposed method, Degree Matrix Comparison (DMC), is a method that fits the task. Unsupervised methods with a focus on graph geometry include: REGAL [5], FINAL [6], Klau [7], and IsoRank [8]. For a more comprehensive view of past unsupervised alignment methods, see [9]. DMC was constructed with two core ideas in mind: degree is the most accurate and direct description of local structure; moreover, in representing geometric structure, it makes sense to include both local and global information, like in [10, 11], especially in compact matrix form (adjacency matrices and Laplacian matrices).
We propose foregoing the need to know exactly which nodes are connected, like in an adjacency matrix, and record accurately the degree of a node and its neighbors’ degrees in a matrix. We show that DMC works well for aligning real world graphs with high clustering and link complexity.
2 Problem Set Up
Definition 2.1 (Graph).
A graph is a pair $G=(V,E)$, where elements $v\in V$ are called vertices (or nodes), and $E$ is a set consisting of unordered (for our purposes only) pairs $\{v_{1},v_{2}\}$ for some but not necessarily all of $v_{1},v_{2}\in V$.
Geometrically, $G$ could represent an object with vertices and edges (links) that connect some of the vertices. The presence of $\{v_{1},v_{2}\}\in E$ indicates there is an edge between the vertices $v_{1},v_{2}$. When we mention graphs in this article, we typically talk about them in a manner that assumes a geometric interpretation. But there are other ways to interpret what we call a “graph”; a most direct example is the adjacency matrix, which is a binary matrix representation of graphs.
Definition 2.2 (Graph Alignment).
If we have $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ for graph alignment, we try to find the optimal mapping $f:V_{1}\rightarrow V_{2}$ to maximize the number of $v_{1}\in C\subseteq V_{1}$ such that $v_{1}=v_{2}=f(v_{1})\in V_{2}$. This process is called Graph Alignment.
To perform graph alignment, we generate a pair of graphs with common structure. We introduce two graph sampling methods–these methods sample size-wise manageable smaller graphs from larger complex networks.
2.1 Random Walk Graph Sampling
One of our methods utilizes random walk on networks to extract subgraphs [3, 12].
Definition 2.3 (Random Walk (on graphs)).
Suppose we start random walk at $v\in V$. We first find the set of neighbors of $v$, which are vertices connected to $v$ by at least one edge, that we denote as $S_{nei}$. Then we randomly pick $v_{n}\in S_{nei}$, and move from $v$ to $v_{n}$. Then we do the same thing for $v_{n}$, and all following nodes iteratively, until called to stop. This process is called random walk.
We take subgraph $G_{s}=(V_{s},E_{s})$, with $n=|V_{s}|$ distinct nodes, of a real world network $G_{r}=(V_{r},E_{r})$ through random walk. This is achieved through the following procedures: we pick random $v_{0}\in V_{r}$, and initiate random walk on $G_{r}$ until we either get one connected component (a graph that cannot be represented as the disjoint union of two graphs) with $n$ distinct nodes, or exhaust a connected component in $G_{r}$ with insufficient nodes, whichever comes first. If the latter is true, we randomly pick a new (different from all previous nodes that we walked on) node $v_{n}\in V_{r}$ and continue random walk until we obtain a sampled graph with the desired $n$ distinct nodes and as few connected components as possible. $V_{s}$ is the set that contains the $n$ distinct nodes from the random walk process just described. $E_{s}\subseteq E_{r}$ inherits edges between sampled nodes in $V_{s}$ from $G_{r}$.
To construct $G_{1}=(V_{1},E_{1})\subseteq G_{s}$ and $G_{2}=(V_{2},E_{2})\subseteq G_{s}$ for graph alignment, we perform random walk on $G_{s}$ to first obtain a set of nodes that serves as the common set of nodes $C=V_{1}\bigcap V_{2}\subset G_{s}$. The number of vertices we want to stop at, for the random walk, is $|C|=np$, where $n$ is like before the number of distinct nodes in $G_{s}$ and $p$ is what we call the “overlap percentage” (with respect to $G_{s}$). We then take arbitrary $H_{1},H_{2}\subseteq(V_{s}\setminus C)$ such that $|H_{1}|=|H_{2}|$ and $V_{s}\setminus C=H_{1}\sqcup H_{2}$ (disjoint union). Let $V_{1}=C\sqcup H_{1}$ and $V_{2}=C\sqcup H_{2}$, then $|V_{1}|=|V_{2}|$. $E_{1}\subseteq E_{s}$ is the inherited set of edges between nodes in $V_{1}$ from $E_{s}$, likewise for $E_{2}$.
Figure 1: Graph Sampling with Random Walk
Figure 1 is a demonstration for what we are trying to take off the entire graph. There are three green circles containing different regions of the graph; the two larger ones contain the same number of nodes, and the smallest circle contains the common region between the two possible sampled graphs through random walk. Note that this is not the only sampling result possible.
2.2 Edge Deletion Graph Sampling
The second graph sampling method is learned from [5]. Suppose $G_{s}$ is the same graph as in section 2.1, then let $G_{1}=G_{s}=(V_{1},E_{1})$, and we delete edges of $G_{1}$ with probability $p_{d}$ to obtain another graph, which is $G_{2}$. If $G_{2}=(V_{2},E_{2})$, then $V_{2}=V_{1}$, and $E_{2}\subseteq E_{1}$.
Figure 2: Graph Sampling with Edge Deletion
Figure 2 is a simple demonstration of the edge deletion graph sampling method, where two edges from the left graph were randomly deleted to obtain the right graph.
3 Proposed Graph Alignment Methods
3.1 Degree Matrix Comparison
Our main proposed method requires creating two degree matrices and comparing them. We denote $\text{deg}(v)$ as the degree of a node $v$, where degree is simply the number of edges connected to a vertex, or more formally, the number of unordered pairs in the edge set $E$ that contains the vertex.
Definition 3.1 (Degree Matrices).
Consider a pair of graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$. By our problem set up, $N=|V_{1}|=|V_{2}|$. Let $V_{1}=\{v_{i}\}_{i=1}^{N}$ and $V_{2}=\{w_{j}\}_{j=1}^{N}$. Let $m_{1}=\text{max}(\{\text{deg}(v_{i})\}_{i=1}^{N})$ and $m_{2}=\text{max}(\{\text{deg}(w_{j})\}_{j=1}^{N})$, and $m=\text{max}(m_{1},m_{2})$. Then the degree matrix $M_{1}$ for $G_{1}$ is a matrix of dimension $(N,m)$, where each row left-aligns, for a unique $v_{1}\in V_{1}$, its neighbors’ degrees in ascending order, with zero-padding to the right. We define $M_{2}$ similarly.
The order of rows in $M_{1}$ and $M_{2}$ is arbitrary. After obtaining $M_{1}$ and $M_{2}$, we use the Munkres assignment algorithm to minimize cost between rows of the matrices, producing a map between rows of $M_{1}$ and $M_{2}$. We migrate this assignment directly to the assignment $f:V_{1}\rightarrow V_{2}$ between nodes. We call this entire process the Degree Matrix Comparison (DMC).
For hands-on examples to demonstrate the concept of a degree matrix and the DMC process, see Appendix A. If the reader is not familiar with the Munkres assignment algorithm, see Appendix B for details. In practice (implementing in code), the Munkres assignment algorithm can be replaced by a computationally more efficient linear_sum_assignment function from scipy.optimize in Python which yields the same assignment as Munkres [13]. But in terms of comprehension, the Munkres assignment algorithm is more straightforward.
The complexity of DMC (either theoretically with Munkres or computationally with linear_sum_assignment) is at most $O(N^{3})$ for a pair of graphs that each has $N$ nodes [14, 15]. When $m<<N$, the complexity is $O(N^{2}\cdot m)<<O(N^{3})$.
3.2 Variation: Greedy DMC
The complexity of DMC can be reduced through replacing direct Munkres assignment with a combination of the auction algorithm and Munkres assignment algorithm, but accuracy is compromised [16]. The auction algorithm, with $O(N^{2})$ complexity and assignment threshold $\epsilon$, serves as a preparation step to reduce reliance on the Munkres assignment algorithm. We call this reduced version the Greedy DMC.
For details about the auction algorithm, one could refer to Appendix C. Here we bring to attention the case where $\epsilon=0$. While this is a special case of the Greedy DMC algorithms with different $\epsilon$ thresholds, it is not “Greedy” in any sense. In fact, this is an algorithm that first finds exact matches (same number of first neighbors and its neighbors’ degrees follow exact same distribution) between rows in $M_{1}$ and $M_{2}$, then performs the Munkres assignment algorithm on other rows. As can be seen in section 5 with the results, Greedy DMC’s $\epsilon=0,10$ cases are very effective algorithms like DMC, and are worth studying in the future.
4 Theoretical Motivations
One theoretical motivation for the method is that degree matrices remain invariant in response to arbitrary node labeling. Moreover, a degree matrix ($N\times m$) encodes information more compactly than an adjacency matrix ($N\times N$). Additionally, combining local (degrees) and global (assignment) perspectives is important because the Friendship paradox shows how local observations can be completely shifted when we examine the same neighborhood globally [17, 18].
The Friendship paradox, in general terms, is the observation that friends of a person typically have more friends than that person. There are many ways to mathematically formulate the problem, and this leads to a whole other field of study; we will not expand on the details here.
Intuitively, our method works well for heterogeneous graphs for the following reason. Suppose there is a graph that has nodes with degrees that are almost all unique. (We phrase it this way with the following consideration in mind: If we allow at most one edge between every pair of vertices on a graph $G_{example}$ with $n_{example}$ vertices and require at least degree one for all nodes, then the maximum degree is at most $n_{example}-1$ for every node, and in this case it is impossible to create a bijection between possible degrees and the nodes.) Then most nodes have row vectors that are relatively unique and distinguishable, compared to, for example, a large set of degree one nodes all connected to a common center.
5 Experimental Results and Analysis
5.1 Real World Networks
Here is a table that concludes the wide range of datasets we worked with.
COSNET [19] Flickr; Last.fm; Myspace
SNAP [20] Facebook; YouTube; CA, PA Roadmaps; PPI
NetworkRepository [21] All other networks.
Method Correct Percentage Approx. Complexity
DMC $96.09\%$ $\sim O(n^{2})-O(n^{3})$
Greedy DMC ($\epsilon=100$) $50.46\%$ $\sim O(n^{2})-O(n^{3})$
REGAL [5] $80\%<p<90\%$ $\sim O(n)-O(n^{2})$
FINAL [6] $35\%<p<45\%$ $\sim O(n^{2})$
Klau [7] $20\%<p<30\%$ $\sim O(n^{5})$
IsoRank [8] $5\%<p<15\%$ $\sim O(n^{4})$
Two-Step $4.42\%$ $\sim O(n^{2})-O(n^{3})$
Euc Dist. $0.15\%$ (near zero) $\sim O(n^{2})-O(n^{3})$
Table 1: Comparison with baselines on PPI network [4, 5, 6, 8]. Deletion probability $p_{d}=0.01$. “Two-Step” uses degrees for alignment, and refines it through average degrees of neighbors. “Euc Dist.” aligns two graphs by minimizing distances between embeddings. DMC has highest accuracy, but larger complexity compared to [5, 6].
(a) Social Networks
(b) Roadmaps
Figure 3: We use the random walk graph sampling method for these experiments.
Networks Mean Variance Total Nodes ($N$)
Flickr 52.36 5330.29 5000
Last.fm 26.03 2164.02 5000
Myspace 7.59 113.32 5000
Facebook 85.06 4040.27 700
YouTube 14.52 960.27 5000
CA Roadmap 2.55 0.96 1000
PA Roadmap 2.88 0.78 1000
Table 2: Means and variances of degree distributions of $G_{s}$. The best performing Facebook and YouTube have large variances in their degree distributions. Comparing Flickr, Last.fm, and Myspace, Flickr and Last.fm are much more heterogeneous. However, there is no strict correlation between variance and performance, as can be seen from California and Pennsylvania roadmaps.
Networks Percent Correct Density Avg. Clustering
bio-grid-fission-yeast .9932 .0123 .1874
bn-mouse-retina-1 .9900 .9983 1.5539
bio-ce-gn .9894 .0218 .1839
bn-fly-drosophila-medulla-1 .9378 .0211 3.8729
bio-grid-human .8440 .0014 .0800
bn-mouse-kasthuri .6156 .0032 0
bio-yeast .5822 .0018 .0708
Table 3: Comprehensive table of experiments on biological networks, ordered by DMC’s performance on $p=90\%$ overlap subgraphs. Density highlight threshold: .01, average clustering coefficient highlight threshold: .1. DMC works better on graphs with higher densities and clusterings.
(a) Performance Decay Plot
$\epsilon$ Correct Percentage
0 $96.58\%$
10 $96.32\%$
100 $50.46\%$
1000 $0.18\%$
(b) Greedy DMC Performance
Figure 4: Experiments on PPI, with random deletion sampling method.
5.2 Graphs from Generation Models
5.2.1 Different Graph Generation Models
In this section, we introduce the Erdos-Renyi, Barabasi-Albert, and Chung-Lu models that were derived to generate complex networks, potentially to model real world network behaviors.
The Erdos-Renyi model is the most basic random graph generation model. One common way to define this is to fix the number of vertices, say $n_{er}$, then connect an edge between every pair of vertices with some probability $p_{er}$. The resulting graph is typically denoted as $G(n_{er},p_{er})$ [22]. One interesting property of these graphs is that their degree distributions are binomial, as can be be seen by the following calculation:
$\textbf{P}(\text{deg}(v)=k)=\binom{n-1}{k}p^{k}(1-p)^{n-1-k}.$ (1)
The probability considers the possibility of having degree of $k$ (being connected to $k$ edges) when the largest possible degree is $n-1$, given the total number of nodes is $n$. $p$ in (1) is the same $p_{er}$ for generating the random graph.
The Barabasi-Albert model considers properties observed on real world networks, especially the preferential attachment phenomena [23]. Instead of attaching new edges with uniform probability, it is possible to attach new edges with weighted probabilities. Suppose we already have $m$ nodes with some connections, then for the $(m+1)^{\text{th}}$ node, the probability that it will be attached to a previous node $v_{i}$ is
$\textbf{P}(v_{i})=\dfrac{\text{deg}(v_{i})}{\Sigma_{i=1}^{m}\text{deg}(v_{i})}$ (2)
We can observe that the probabilities of attaching the new node to any previous node sums up to one, forming a probability distribution; moreover, higher degree nodes attract more new nodes. This usually leads to what we call hubs, which are local centers that are attached to many low degree nodes.
Now we introduce the Chung-Lu model. It also attempts to capture real world network properties, but instead of focusing on preferential attachment or community behaviors, they attempt to approximate real world networks’ degree distributions [24]. Suppose we have a specific degree distribution in mind and we expect $m$ nodes in total, then the attachment probability between each node $v_{i}$ and another node $v_{j}$ is as follows:
$\textbf{P}(ij)=\dfrac{\text{deg}(v_{i})\text{deg}(v_{j})}{\Sigma_{k=1}^{m}\text{deg}(v_{k})}$ (3)
Moreover, the expected degree at each node $v_{i}$ is
$\Sigma_{k=1}^{m}\textbf{P}(ik)=\Sigma_{k=1}^{m}\dfrac{\text{deg}(v_{i})\text{deg}(v_{k})}{\Sigma_{k=1}^{m}\text{deg}(v_{k})}=\text{deg}(v_{i})$ (4)
which theoretically guarantees that the Chung-Lu model will generate a synthetic graph with a degree distribution that is close to our initial distribution in mind.
5.2.2 Discussion on Model Approximations to Real World Networks
Through examining existent literature, we summarize a set of (non-comprehensive) key words that are properties observed from real world networks: hierarchical [25], fractal [26], copying behavior [23], preferential attachment [23], small world [27], clustering [28], power-law degree distribution [23]. These properties are of course not enough to capture the entirety of real world networks, which makes building new models to generate complex networks interesting.
5.2.3 Applying DMC to Generated Graphs
Definition 5.1 (k-partite Graph).
A k-partite Graph is a graph with nodes that can be split into $k$ independent sets, where nodes within each set are never connected.
We applied DMC to graphs generated through the three models (with some slight variations). We examine the performance of DMC on the graphs with respect to different partite numbers $k$.
For our Erdos-Renyi experiments, we fix total number of nodes to be $n_{er}=100$, then split the nodes into $k$ partite groups; the partition of nodes is arbitrary. After that, we connect edges between different partite groups with probability of $p_{er}=0.5$. Then we use the random walk graph sampling method to obtain two graphs with 90% overlap, and perform DMC. We run the same experiment for 20 times, then take the mean performance.
For our Barabasi-Albert experiments, we do a similar thing. Every time we add a new node, we connect it to five different previous nodes (add 5 edges). For our Chung-Lu experiments, we use a power-law distribution as our initial degree distribution, where the power-law exponent was chosen to be $\gamma=2.5$, and we set a maximum degree limit of 20.
Figure 5: Discovery of negative linear relationship.
Figure 6: BA model has much better performance than ER. (It is more heterogeneous.)
Figure 7: CL model with initial power-law degree distribution.
Real world networks are more complex than we can imagine, hence the difficulty of capturing them with mathematical language. However, it is interesting the simple DMC works better on the more complex real world networks than synthetically generated graphs. This shows that modeling complex networks is no easy task, and understanding the complexity will potentially bring us rich information.
6 Conclusion and Future Work
We proposed the DMC and Greedy DMC, showed that they were theoretically motivated, and analyzed experimental results. DMC’s accuracy on classic PPI networks proves to be higher than carefully chosen baselines, and it works best for heterogeneous graphs. Future works could examine special cases of Greedy DMC, as pointed out before. We also provided a brief overview of synthetic graph generation models and a short discussion of the difficulty of modeling complex networks.
7 Acknowledgements
We thank Allison Gunby-Mann for introducing Ashley to the graph alignment problem and bringing the random walk sampling method to attention.
Appendix A Examples to Demonstrate DMC
A.1 Degree Matrix Formulation
Let us examine the graph (assumed to be a local picture taken from a much larger graph) in Figure 8 with circled node as the origin. Then the boxed nodes are the first neighbors of the origin. Starting from the first neighbor in the upper left corner and going counter clock-wise, the degrees of the five first neighbors are $3,4,3,5,3$. Suppose (arbitrarily) $m$ in this case is $10$, then the row vector representation for the circled node would be $[3,3,3,4,5,0,0,0,0,0]$ in the degree matrix for the larger graph. We do this for every node in the larger graph which this smaller piece is taken from to build a degree matrix.
Figure 8: Example Graph to Demonstrate Degree Matrix Formulation
A.2 DMC Demonstration
Now consider the two graphs in Figure 9 as a pair that we are to align as a simple example.
(a) $F_{1}$
(b) $F_{2}$
Figure 9: Example Graphs to Demonstrate DMC
Let the graph in Figure 9 (a) be $F_{1}=(V_{F_{1}},E_{F_{1}})$ and the graph in Figure 9 (b) be $F_{2}=(V_{F_{2}},E_{F_{2}})$. The node set $V_{l}$ of the largest overlapping subgraph consists of all five nodes, or in mathematical notation: $V_{l}=V_{F_{1}}=V_{F_{2}}$. However, the graphs are not isomorphic, so the edge set $E_{l}$ of the overlapping subgraph is a strict subset of $E_{F_{1}},E_{F_{2}}$, which in mathematical notation is: $E_{l}\subsetneq E_{F_{1}},E_{F_{2}}$. Let the edge between node 3 and node 4 in $F_{2}$ be $w$. Then the overlapping subgraph takes the shape of $F_{2}\backslash\{w\}$. The degree matrices for $F_{1}$ and $F_{2}$ are
[…]
Then we use the Munkres assignment algorithm to minimize the cost of row alignment between the two matrices above. The results indicate that the algorithm aligns nodes $1,2,3,4,5$ of $F_{1}$ to nodes $2,5,1,3,4$ of $F_{2}$ (one could check understanding of the concepts with this result). This is not an optimal result, but it is expected since the graphs in this example have low heterogeneity. This example is purely for the purpose of demonstrating the method.
Appendix B Munkres Assignment Algorithm
Algorithm 1 Munkres Assignment Algorithm
1:  Input: A cost matrix $C$ of size $n\times n$.
2:  Output: An optimal assignment and the total cost.
3:  Step 1: Subtract the row minimum.
4:  for each row $i$ in $C$ do
5:     Subtract the minimum value of row $i$ from all elements in row $i$.
6:  end for
7:  Step 2: Subtract the column minimum.
8:  for each column $j$ in $C$ do
9:     Subtract the minimum value of column $j$ from all elements in column $j$.
10:  end for
11:  Step 3: Cover all zeros with a minimum number of lines.
12:  while not all zeros are covered do
13:     Identify rows and columns containing uncovered zeros.
14:     Cover all zeros using the minimum number of horizontal and vertical lines.
15:  end while
16:  Step 4: Check the number of lines.
17:  if the number of lines is equal to $n$ then
18:     An optimal assignment is possible.
19:     Go to Step 6.
20:  else
21:     Go to Step 5.
22:  end if
23:  Step 5: Adjust the matrix.
24:  Find the smallest uncovered value.
25:  Subtract this value from all uncovered elements.
26:  Add this value to elements at intersections of covering lines.
27:  Return to Step 3.
28:  Step 6: Make the assignments.
29:  Select zeros such that no two are in the same row or column.
30:  Return the assignment and compute the total cost.
Appendix C Auction Algorithm
Algorithm 2 Auction Algorithm with Munkres for Unassigned Items
0:  Profit matrix $P$ of size $n\times m$ (agents $\times$ tasks), threshold $\epsilon>0$.
0:  Assignment of agents to tasks.
1:  Initialize $prices[j]\leftarrow 0,\;\forall j\in\{1,\ldots,m\}$.
2:  Initialize $assignments[i]\leftarrow\text{unassigned},\;\forall i\in\{1,\ldots,n\}$.
3:  while any agent $i$ is unassigned do
4:     for each unassigned agent $i$ do
5:        Compute net values $v[j]\leftarrow P[i,j]-prices[j],\;\forall j$.
6:        Identify the best task: $j_{1}\leftarrow\arg\max_{j}v[j]$.
7:        Identify the second-best task: $j_{2}\leftarrow\arg\max_{j\neq j_{1}}v[j]$.
8:        Compute bid increment: $\Delta\leftarrow v[j_{1}]-v[j_{2}]+\epsilon$.
9:        Update the price: $prices[j_{1}]\leftarrow prices[j_{1}]+\Delta$.
10:        Assign task $j_{1}$ to agent $i$.
11:        if task $j_{1}$ was previously assigned to agent $k$ then
12:           Unassign agent $k$.
13:        end if
14:     end for
15:  end while
16:  if any tasks remain unassigned then
17:     Assign remaining tasks using the Munkres assignment algorithm.
18:  end if
19:  Return: Final assignment of agents to tasks.